Accountancy

Answer parts A,B,C,D and E to receive full credit for #1:

A. Becks company includes one coupon in each bag of cat food it sells. In return for 15 coupons and 25 cents, customers recieve a leash. the leashes cost the compny $4.50 each. Beck estimates that 40% of the coupons will be redeemed. Data for 2019 are as follows:


2019
bags of cat food sold1,650,000
coupons redeemed225,000

Beck does not make any coupon expense entries at the time of sale. rather, it recognizes coupon expense at the time of actual redemptions, and makes an adjuting entry at the end of the year.

how much coupon (premium) expense should Beck record for 2019?

1. $63,750

2. $187,000

3. $198,000

4. $225,000

B. how much is Becks estimated liability for coupons (premiums) at December 31,2019?

1. $0

2. $123,250

3. $187,000

4. $225,000

C. The total payroll of a company for the month of october 2019 was $280,000. This amount is entirely subject to F.I.C.A (6.2%), medicare (1.45%), federal (0.8%), and state (3%) unemployment taxes. Also, $63,000 of federal income taxes and $1,100 of union dues were withheld. The journal entry to record the wages paid and employee payroll deductions:

Account Titledebitcredit


















1. reduces total assets by $85,520

2. reduces net income by $194,480

3. reduces total liabilities by $85,520

4. reduces total assets by $194,480

D: how much employer payroll tax expense should Beck recognize for October 2019?

1. $21,420

2. $22,520

3. $29,820

4. $32,060

E. which of the following statements is/ are correct?

(i)  amortization of discount on bonds payable increases the carrying value of bonds payable.

(ii) Amortization of premium on bonds payable increases the carrying value of bonds payable

(iii)  the entry to record a coupon payment made by a bond issued at premium has no effect on total assets

1. i only

2. ii only

3. i and iii

4. ii and iii
Answer #1

A.
Premium Expense = (1650000 x 40%) / 15 x ($4.50-0.25) = $187,000
Answer is 2. $187,000

B.
Estimated Liability for Coupons = (1650000 x 40%-225000) / 15 x ($4.50-0.25) = $123250
Answer is 2. $123,250

C.
Answer is 4. reduces total assets by $194,480

Account Titles Debit Credit
Salaries Expense $          280,000
      FICA Tax Payable $            17,360
      Medicare Tax Payable $              4,060
      Federal Income tax Payable $            63,000
      Union Dues Payable $              1,100
      Cash $          194,480

D.
Payroll Tax Expense = $17360 + 4060 + 280000 x 3.80% = $32060
Answer is 4. $32,060

E.
Answer is 1. i only

If you have any query, kindly comment with your query and please mark thumbs up.